the difference of the squares of two consecutive integers is 24 what are the integers
Answer by rwm(914)

(x+1)^2-x^2=24
x^2+2x+1-x^2=24
2x+1=24
2x=23
x=23/2
Something is wrong with your problem.
It probably should read "consecutive odd integers"
(x+2)^2-x^2=24
x^2+4x+4-x^2=24
4x+4=24
4x=20
x=5
x+2=7
check
49-25=24
24=24
ok